If totally 25 workers working for 7 hours a day completing a particular piece of work in 28 days,in howmany days 20 workers working for 6 hours a day wil complete same work?
Answer by ankor@dixie-net.com(22740)

If totally 25 workers working for 7 hours a day completing a particular piece
of work in 28 days, in how many days 20 workers working for 6 hours a day will complete same work?
:
Find number of man-hours required to complete the job.
25*7*28 = 4900 man-hrs
:
let d = number of days required by 20 workers, working 6 hr/day
20*6*d = 4900
120d = 4900
d = 4900/120
d = 40.833 days which is: 40 days + .833(6) = 5 hrs
:
Check this
40*6*20 = 4800 man-hrs
20 * 5 = 100 man-hrs
----------------------
total 4900 man-hrs